# SLoRD: Structural Low-Rank Descriptors for Shape Consistency in Vertebrae Segmentation

### v1

@@ 3.3 Ablation Studies
We also carried out the ablation study on the plug-and-play property of SLoRD. For MedNeXt [22], the introduction of SLoRD brings remarkable gains on the segmentation performance of the public and hidden test datasets, with $2.43\%$, $1.96\%$ Dice score increases and $4.45mm$, $4.82mm$ HD decreases. Besides MedNeXt, our proposed SLoRD will boost segmentation networks including Verteformer [7], Spine-Transformer [15], [16] as shown in Table 2. Specifically, for Verteformer [7] which is aimed at the task of vertebrae segmentation, substantial improvements have been achieved on the average Dice score and HD ($2.59\%\uparrow$, $3.09mm\downarrow$). Detailed qualitative results are provided in Figure 8 to demonstrate the efficacy of SLoRD in promoting segmentation consistency when plugged into single-stage, two-stage, and multi-stage methods. As long as there is a satisfactory performance for the top two vertebrae, even with moderate inconsistency, SLoRD can significantly boost segmentation results.
@@ 4 Conclusion
In this paper, we address the challenge of intra-vertebrae segmentation inconsistency from the perspective of instance segmentation. Specifically, we propose a contour-based network to refine coarse predictions, based on the linear combination of structured vertebral contour descriptors. Besides, the proposed contour loss imposes consistency constraints, facilitating contour points close to vertebral boundaries. Detailed quantitative and qualitative results on VerSe 2019 have demonstrated the efficacy of our devised framework.
Acknowledgments
This work is supported in part by National Key R&D Program of China (2019YFB1311503), the Shanghai Sailing Program (20YF1420800), the Shanghai Health and Family Planning Commission (202240110), and National Key R&D Program of China (62373243).

### v2

@@ 4.3.3 Plug-and-play
We also carried out the ablation study on the plug-and-play property of SLoRD on VerSe 2019 and 2020. Besides MedNeXt [32], the proposed SLoRD will boost segmentation results achieved by Verteformer [45], Spine-Transformer [38], Payer’s method [31] as shown in Table 6. Specifically, for Verteformer on VerSe 2020, substantial improvements have been achieved on the average Dice score and HD ($2.65\%\uparrow$, $2.16mm\downarrow$). Also, SLoRD can repair segmentation results by Payer’s approach, with $2.01\%\uparrow$ average Dice score and $0.48mm\downarrow$ average HD. Qualitative results are provided in Fig.11 to demonstrate the efficacy of SLoRD in promoting segmentation consistency when plugged into single-stage, two-stage, and multi-stage methods.
Figure 11: Improvement of our proposed SLoRD when applied to coarse predictions from other segmentation networks including the single-stage Verteformer [45], two-stage Spine-Transformer [38], and multi-stage Payer’s method [31].
4.3.4 Comparison with PCA-based shape model
As we mentioned in Section 3.1, SVD shows the best low-rank approximation property and can mine more robust shape descriptors compared with PCA. Here a straight comparison is implemented on the hidden test dataset of VerSe 2020. As illustrated in Table 9, our method outperforms DeepSSM significantly, with a $3.92\%$ higher Dice score and a $1.73mm$ lower Hausdorff distance. Also, since different vertebrae bear various numbers of contour points, it is tricky to carry out the uniform sampling process while delineating the original contour. In contrast, our method achieves a $1.117\downarrow$ ASD value, showing more satisfactory contour reconstruction results based on the structural encoding way of contour representations.
@@ 5 Conclusion
In this paper, to address the challenge of intra-vertebrae segmentation inconsistency, we devise SLoRD by jointly imposing voxel consistency and contour precision regularizations. Specifically, we propose a structural contour representation way based on the spherical coordinate system. Besides, basic contour descriptors and shape constraints facilitate regressed contour points close to vertebral surfaces. Detailed quantitative and qualitative results on VerSe 2019 and 2020 demonstrate the superiority of our method compared with other single-stage, two-stage and multi-stage models. Moreover, SLoRD serves as a plug-and-play framework to repair the segmentation in- consistency in coarse predictions from other methods.